initially at human body temperature and 725mmHg, a 456mL sample of gas is cooled to -25.0 celsius while having pressure changed what is the volume of the sample
1
Expert's answer
2020-12-14T14:45:02-0500

Gas Equation, "PV = n R T" holds good in all situations.

Given, Pressure unchanged and no. of particles fixed,

P1V1= n R T1 and P2V2= n R T2

Dividing them, we get "\frac{V1}{V2} = \frac{T1}{T2}"

we need to find V2,

Therefore V2 = "\frac{V1T2}{T1}"

= 456 * 248/310

= 364.8 mL